OpenStudy (anonymous):

hat is the slope of the line that passes through (3, –7) and (–1, 1)

OpenStudy (anonymous):

The slope of any line through two points (X1, Y1) and (X2, Y2) is calculated as follows: Subtract the two Xs, Subtract the two Ys Slope = Y-result / X-result. In your case: X-result = 3 - (-1) = 4 Y-result = -7 - 1 = -8 Slope = Y-result / X-result = -1 / 4 = -2